OpenStudy (anonymous):

find the zeros of f(x)=3x^3-x^2-2x ,f(x)=X^3-25x ,f(x)=7(x-3)^2(x+5)^4

hero (hero):

1. Factor 2. Set f(x) = 0 3. Solve 0 = x(x^2 - 25) 0 = x(x + 5)(x - 5) 0 = x 0 = x + 5 0 = x - 5 0 = 7(x - 3)^2(x + 5)^4 0 = (x - 3)^2(x + 5)^4 0 = (x - 3)^2 0 = (x + 5)^4 0 = x - 3 0 = x + 5

hero (hero):

Basically, I helped you with the last two

hero (hero):

By the way, just in case you were wondering The square root of zero is zero The quad root of zero is zero

OpenStudy (anonymous):

thanks a lot you are really my hero :D

hero (hero):

For the first one: 0 = x(3x^2 - x - 2) 0 = x(3x^2 - 3x + 2x - 2) 0 = x(3x(x - 1)+2(x - 1)) 0 = x((x - 1)(3x + 2)) 0 = x(x - 1)(3x + 2)

hero (hero):

Once you have completely factored, you can set each factor equal to zero, then solve for x.

OpenStudy (anonymous):

on the second one what happen to the 7?

hero (hero):

We have an equation. In an equation if you have something like 0 = 7x You can divide both sides by 7 to get x = 0 If you have something like 0 = 7(x + 4)(x - 3) It is still legal to divide both sides by 7 to get 0 = (x + 4)(x - 3)

hero (hero):

Remember, 0/7 = 0

hero (hero):

You can divide zero by any number other than zero and you will get zero. What you cannot do is divide by zero.

hero (hero):

I hope that doesn't confuse you

OpenStudy (anonymous):

ohhh ok thanks and nop im not confused, yet

hero (hero):

Okay good.

OpenStudy (anonymous):

thanks a lot you are truly a hero